# Mixed birth-death and death-birth updating in structured populations

### v1

@@ 4 Conclusion
In this paper, we have studied mixed dB and Bd updating for neutral evolution and for constant selection. At each time step we use dB updating with probability $\delta$ and dB updating with probability $1-\delta$. We have derived the following results that appear both in the main text and the appendices in full.
The fixation probability as a function of $\delta$ can be increasing, decreasing, or possibly non-monotonic in neutral evolution on an undirected unweighted graph. Despite this, we find that the fixation probability is $1/N$ on any such graph when $\delta=1/2$. Additionally, it cannot be the case that the fixation probability is greater than $1/N$ for both pure dB and pure Bd (see Appendix A).
Similar to the fixation probability, the fixation time as a function of $\delta$ can be increasing, decreasing, or possibly non-monotonic in neutral evolution on an undirected unweighted graph. We prove that the expected (conditional) fixation time on such graphs is at most some polynomial in $N$ when $\delta=1/2$ and $r>1$ (see Appendix E). We also prove that for $\lambda$-almost regular graphs, the expected fixation time is at most a polynomial in $N$ for any $\delta$ and when $r\geq\lambda^{2}>1$ (see Appendix G). When $r>1$, we provide an efficient estimation algorithm for any $\delta$ for the fixation probability on almost all unweighted undirected graphs (see Appendix H).
For weighted and possibly directed graphs that are a uniform circulation, the fixation probability in neutral evolution is $1/N$, regardless of $\delta$ (see Appendix B).
For constant selection on undirected cycles and stars we provide tractable analytical formulas for all $\delta$ (see Appendix I and Appendix J, respectively). For neutral evolution, we give an explicit formula for the fixation probability on a bidegreed graph. Further, we prove that with a deterministic initial mutant location, the fixation probability in neutral evolution on a bidegreed graph is the most sensitive at pure Bd or pure dB (see Appendix K).
There are many future directions of this work. We provide three promising directions. Firstly, one could classify the graphs that have monotone fixation probability in $\delta$, even in neutral evolution. Secondly, one could find for each $\delta$, which graphs maximize fixation probability when $r>1$ under uniform initialization. Finally, one could estimate a realistic $\delta$ from data, by fitting the $\delta$-mixed updating model to empirical fixation probabilities in, for example, microbial populations, cancer cell lines, or laboratory evolution experiments.
Data Accessibility. Our code is available as a GitHub repository at https://github.com/davidb2/mixed-updating. For graph exploration, we utilized the computer program nauty [47].
Acknowledgements. D.A.B. was supported by a Harvard Graduate School of Arts and Sciences Prize Fellowship. Y.H. and M.M. were supported by NSF grant CNS-2107078.
@@
Appendix A Proof that it is not possible for both pure processes to simultaneously have fixation probabilities higher than the unbiased process
Theorem A.1.
Let $G=(V,E)$ be an undirected, unweighted, connected graph. Let $u\in V$. Then it must be the case that
@@
Appendix B Proof of Theorem 3.2
Lemma B.1 (Additivity in neutral evolution).
$\mathrm{fp}_{r=1}^{\delta}(G,S)=\mathrm{fp}_{r=1}^{\delta}(G,T)+\mathrm{fp}_{r=1}^{\delta}(G,S\setminus T).$ (14)
Proof.
Consider the variant of the process where every vertex starts with its own mutant, all with fitness $1$. The process will evolve as usual until a single color takes over the whole graph. A mutant currently occupying subset $S$ fixates if and only if one of the vertices in the subset takes over in the altered process, which in turn equals the fixation probability with only that vertex by viewing all other colors as wild-types. ∎
Definition B.1.
Let $G=(V,E)$ be a directed graph with vertices labelled $V:=\{1,\ldots,N\}$ and edge weights $w\colon V^{2}\to\mathbb{R}_{\geq 0}$. Note that $w_{uv}>0$ if and only if $u\to v\in E$. Let $w_{uv}^{\prime}:=\frac{w_{uv}}{\sum_{v^{\prime}\in V}w_{uv^{\prime}}}$ so that $\sum_{v\in V}w_{uv}^{\prime}=1$. Let $w_{uv}^{\prime\prime}:=\frac{w_{uv}}{\sum_{u^{\prime}\in V}w_{u^{\prime}v}}$ so that $\sum_{u\in V}w_{uv}^{\prime\prime}=1$. The temperature of a vertex $v\in V$ is defined as the probability death occurs at $v$ in the next time step when $r=1$,
$T^{\delta}_{v}:=\delta T^{\textrm{dB}}_{v}+(1-\delta)T^{\textrm{Bd}}_{v}$ (15)
where
$T^{\textrm{Bd}}_{v}:=\frac{1}{N}\sum_{u\in V}w_{uv}^{\prime}\quad\text{and}\quad T^{\textrm{dB}}_{v}:=1/N.$ (16)
Note that for any $\delta\in[0,1]$, we have that for all $u,v\in V$ we have $T^{\delta}_{u}=T^{\delta}_{v}\iff T^{\textrm{Bd}}_{u}=T^{\textrm{Bd}}_{v}$. For this reason, we write $T_{v}:=T^{\textrm{Bd}}_{v}$.
Definition B.2.
A weighted, directed graph $G=(V,E)$ is called isothermal if $T_{u}=T_{v}$ for every pair $(u,v)\in V^{2}$.
Complete (unweighted) graphs are a special type of isothermal graphs. When the population is modeled as a complete graph with self loops, the fixation probability and time are invariant to $\delta$. This is because a Bd update and a dB update have the same dynamics/underlying Markov chain for a complete graph.

### v2

@@ 4 Conclusion
In this paper, we have studied mixed dB and Bd updating for neutral evolution and for constant selection. At each time step we use dB updating with probability $\delta$ and Bd updating with probability $1-\delta$. We have derived the following results that appear both in the main text and the appendices in full.
The fixation probability as a function of $\delta$ can be increasing, decreasing, or possibly non-monotonic in neutral evolution on an undirected unweighted graph. Despite this, we find that the fixation probability is $1/N$ on any such graph when $\delta=1/2$. Additionally, it cannot be the case that the fixation probability is greater than $1/N$ for both pure dB and pure Bd (see Appendix D).
We give an explicit formula for the fixation probability on a bidegreed graph for any $\delta$. Further, we prove that with a deterministic initial mutant location, the fixation probability in neutral evolution on a bidegreed graph is the most sensitive at pure Bd or pure dB (see Appendix E.1).
Similar to the fixation probability, the fixation time as a function of $\delta$ can be increasing, decreasing, or possibly non-monotonic in neutral evolution on an undirected unweighted graph. We prove that the expected (conditional) fixation time on such graphs is at most some polynomial in $N$ when $\delta=1/2$ (see Appendix F.1).
For constant selection with $r>1$, we prove that when $\delta=1/2$ the fixation time is at most $N^{5}\cdot r/(r-1)$ on any undirected unweighted graph (see Section 9 in the SI). We also prove that for graphs with $r>(D/d)^{2}$, the expected fixation time is at most $N^{7}\cdot r/(r-1)$ for any $\delta$ (see Section 12 in the SI). When $r>(D/d)^{2}$, we provide an efficient estimation algorithm for any $\delta$ for the fixation probability (see Section 15 in the SI).
For constant selection on undirected cycles and stars, we provide tractable analytical formulas for all $\delta$ (see Section 10 and Section 11 in the SI, respectively).
While fixation probabilities and times can have strange behaviors with respect to $\delta$, we have shown that unbiased mixing and nearly-regular graphs lead to well-behaved quantities.
There are many future directions of this work. We provide two promising directions. Firstly, one could classify the graphs that have monotone fixation probability in $\delta$, even in neutral evolution. Finally, one could estimate a realistic $\delta$ from data, by fitting the $\delta$-mixed updating model to empirical fixation probabilities in, for example, microbial populations, cancer cell lines, or laboratory evolution experiments.
Data Availability. Our code is available as a GitHub repository at https://github.com/davidb2/mixed-updating. For graph exploration, we utilized the computer program nauty [52]. We also explored dynamics using EvoLudo [53].
Acknowledgments. The authors thank Josef Tkadlec for many helpful suggestions. D.A.B. was supported by a Harvard Graduate School of Arts and Sciences Prize Fellowship. Y.H. and M.M. were supported by NSF grant CNS-2107078.
@@
Appendix A Technical Tools
One main technical tool we will use is the martingale. We will use a potential function on the set $S\subseteq V$. The ultimate goal is to show that variance (i.e. second moment) of the mixed $1/2$-updating process as a function of the potential fluctuates a reasonable amount in expectation. We first recall the optional stopping theorems
Theorem A.1 (Optional Stopping Theorem for martingales, Corallary of Theorem 17.4 of [55]).
Let $k>0$ be some constant. Let $(X_{i})_{i\geq 0}$ be a Markov chain with state space $\mathcal{X}$ such that $\{0,k\}\subseteq\mathcal{X}\subseteq[0,k]$. Let $\tau:=\min\{i:X_{i}\in\{0,k\}\}$. Suppose that $\mathbb{E}[\tau]<\infty$ and $\mathbb{E}[X_{i+1}-X_{i}\mid X_{i}=x]=0$ for all $i\geq 0$ and $x\in\mathcal{X}$. Then $\mathbb{E}[X_{\tau}]=\mathbb{E}[X_{0}]$.
Theorem A.2 (Optional Stopping Theorem for supermartingales, Corallary of Theorem 17.4 of [55]).
Let $k>0$ be some constant. Let $(X_{i})_{i\geq 0}$ be a Markov chain with state space $\mathcal{X}$ such that $\{0,k\}\subseteq\mathcal{X}\subseteq[0,k]$. Let $\tau:=\min\{i:X_{i}\in\{0,k\}\}$. Suppose that $\mathbb{E}[\tau]<\infty$ and $\mathbb{E}[X_{i+1}-X_{i}\mid X_{i}=x]\leq 0$ for all $i\geq 0$ and $x\in\mathcal{X}$. Then $\mathbb{E}[X_{\tau}]\leq\mathbb{E}[X_{0}]$.
Corollary A.1 (Fixation probability of martingale).
Let $k_{2}\geq k_{1}>0$ be some constants. Let $(X_{i})_{i\geq 0}$ be a Markov chain with state space $\mathcal{X}$ such that $\{0,k_{2}\}\subseteq\mathcal{X}\subseteq[0,k_{2}]$. Suppose $\mathbb{E}[X_{0}]=k_{1}$. Let $\tau:=\min\{i:X_{i}\in\{0,k_{2}\}\}$. Suppose that $\mathbb{E}[\tau]<\infty$ and $\mathbb{E}[X_{i+1}-X_{i}\mid X_{i}=x]=0$ for all $i\geq 0$ and $x\in\mathcal{X}$. Then $\mathbb{P}(X_{\tau}=k_{2})=k_{1}/k_{2}$.
Proof.
By Theorem A.2, $\mathbb{P}(X_{\tau}=k_{2})\cdot k_{2}=\mathbb{E}[X_{\tau}]=\mathbb{E}[X_{0}]=k_{1}$. ∎
Lemma A.1 (Absorption time is finite).
We consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $1/2$-updating process where $Y_{i}\subseteq V$. Let $\phi\colon 2^{V}\to\mathbb{R}_{\geq 0}$ and $\tau:=\min\{i:\phi(Y_{i})\in\{0,\phi(V)\}\}$. Then $\mathbb{E}[\tau]<\infty$.
Proof.
Consider a contiguous block of $N$ successive states of the Markov chain $Y_{i},\dots,Y_{i+N-1}$ with $Y_{i}\not\in\{\emptyset,V\}$. Then it must be that $|Y_{i}|\geq 1$. Consider such a contiguous block in which some $u\in S_{i}$ spreads to the whole graph. Each step happens with probability at least $1/N^{2}$, so the entire sequence occurs with probability at least $p=N^{-2N}$. Thus, the expected number of such contiguous blocks until absorption occurs is at most $1/p<\infty$. ∎
Consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $\delta$-updating process where $Y_{i}\subseteq V$. Let $\phi\colon 2^{V}\to\mathbb{R}_{\geq 0}$. We define the following quantities for use in the proofs that follow:
$\psi_{u,v}^{\textrm{Bd}}(S)$ $:=\frac{1}{N}\left(\frac{1}{\deg(u)}(\phi(S\cup\{v\})-\phi(S))-\frac{1}{\deg(v)}(\phi(S)-\phi(S\setminus\{u\}))\right)$ (8)
$\psi_{u,v}^{\textrm{dB}}(S)$ $:=\frac{1}{N}\left(\frac{1}{\deg(v)}(\phi(S\cup\{v\})-\phi(S))-\frac{1}{\deg(u)}(\phi(S)-\phi(S\setminus\{u\}))\right)$ (9)
$\psi_{u,v}^{\delta}(S)$ $:=(1-\delta)\psi_{u,v}^{\textrm{Bd}}(S)+\delta\psi_{u,v}^{\textrm{dB}}(S).$ (10)
Then $\psi_{u,v}^{\delta}(S)=\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]$.
Appendix B Additivity of fixation probability
Lemma B.1 (Additivity in neutral evolution).
$\mathsf{fp}_{r=1}^{\delta}(G,S)=\mathsf{fp}_{r=1}^{\delta}(G,T)+\mathsf{fp}_{r=1}^{\delta}(G,S\setminus T).$ (11)
Proof.
Consider the variant of the mixed $\delta$-updating process where every vertex starts with a unique type. The process will evolve as usual until a single type takes over the whole graph. A mutant currently occupying subset $S$ fixates in the mixed $\delta$-updating process if and only if one of the vertices in the subset takes over in the altered process, which in turn equals the fixation probability with only that vertex by viewing all other types as wild types. ∎
Appendix C Proof of Theorem 3.2
Proof.
We consider the Markov chain $(Y_{i})_{i\geq 0}$ on $\Omega=2^{V}$ of the mixed $1/2$-updating process where $Y_{i}\subseteq V$. Let $S\subseteq V$. We use the potential function $\phi(S)=|S|$. Let $\tau:=\min\{i:\phi(Y_{i})\in\{0,N\}\}$. By Lemma A.1, $\tau$ has finite expectation. Notice that $\psi_{u,v}^{\text{Bd}}(S)=-\psi_{u,v}^{\text{dB}}(S)$, which implies $\psi_{u,v}^{\delta=1/2}(S)=0$. Therefore $\mathbb{E}[\phi(Y_{i+1})-\phi(Y_{i})\mid Y_{i}=S]=\psi_{u,v}^{\delta=1/2}(S)=0$. Then we can apply Corollary A.1 to yield the desired result. ∎
Appendix D Proof that it is not possible for both pure processes to simultaneously have fixation probabilities greater than $1/N$
Lemma D.1.
Let $G=(V,E)$ and $u\in V$. Then it must be the case that
@@ Appendix B Additivity of fixation probability
Appendix E Proof of Theorem 3.5